3. For the reaction
Na2CO3 (s) + 2HCl(g) → NaCl(s) + CO2 (g) + H2O(l)
∆H is -144.1 kJ. What is ∆U.
1
Expert's answer
2020-09-29T06:28:02-0400


We know according to thermodynamics,


∆H = ∆U +∆n.RT____[A]


Where , ∆n= difference in mole no. Of gasious reactant from the number of moles of gaseous product.


R = universal gas constant


T = absolute temperature of the system


This reaction occurs in room temperature.


i.e T=( 25 + 273) K


= 298 K


∆n =number of moles of CO2 - number of moles of HCl


= 1 - 2


= -1


Putting the given values in the above [A] equation become,


-144.1 ×10^3 = ∆U +(-1)×8.314×298


Or,∆U = -144.1×10^3 + (1×8.314×298)


=-141697.254 joul


≈-141.697 kJ

Hence, the ∆U value of the given reaction is = -141.697 kJ